Two identical waves travel in the same direction, each with a wavelength λ = 1 m and speed v = 20 m/s. When the two waves interfere, they form a resultant wave. The angular frequency of the resultant wave is:


40π rad/s


π/10 rad/s


π/5 rad/s


20π rad/s


80π rad/s


1
Expert's answer
2021-05-27T08:43:21-0400

Gives

Wavelength "\lambda=1m"

Speedv=20m/sec

"y=Asin(wt-kx)"

Phase "\phi=(wt-kx)"

Velocity

"V=Awcos(wt-kx)"

Maximum velocity

"wt-kx=0"

cos0°=1

"w=\frac{Kx}{t}"

Put value

"w=\frac{2\pi}{\lambda }\times\frac {x}{t}"

"w=\frac{2\pi}{\lambda }\times v"

Put value

"w=\frac{2\pi}{1}\times20=40\pi Rad /sec"

Option (1) is correct